OpenStudy (anonymous):

Find the limit of t(lnt) as t approaches 0+

OpenStudy (lgbasallote):

\[\lim_{x \rightarrow 0^+} t(\ln t)\] hmm

OpenStudy (lgbasallote):

substitute..

OpenStudy (lgbasallote):

seems like a l'hospital

OpenStudy (turingtest):

\[t=\frac1n\]so as\[t\to0^+\implies n\to\infty\]then yes, l'hospital

OpenStudy (turingtest):

though there is probably a cooler zarkon way to do it :)

OpenStudy (lgbasallote):

if you sub it will be 0 x -infinity so use l'hospital \[-\infty \times \frac{1}{\infty} \implies \frac{-\infty}{\infty}\]

OpenStudy (anonymous):

usual gimmick is to write as \[\frac{\ln(x)}{\frac{1}{x}}\]

OpenStudy (zarkon):

I would just write it as \[\lim_{x \rightarrow 0^+} \frac{\ln t}{1/t}\] and use L'hospitals rule

OpenStudy (turingtest):

oh why do I always forget that^ :S

OpenStudy (lgbasallote):

so turn it into \[\frac{\ln t}{\frac{1}{t}}\]

OpenStudy (zarkon):

change my x to a t

OpenStudy (anonymous):

lol change mine as well

OpenStudy (zarkon):

that is what i get by coping and pasting

OpenStudy (anonymous):

or maybe use \(\xi\)

OpenStudy (lgbasallote):

add \(\lim_{t \rightarrow 0^+}\) to mine

OpenStudy (zarkon):

\[\eta\]

OpenStudy (lgbasallote):

\[\large \lim_{t \rightarrow 0^+} \frac{\frac{d}{dx} \ln t}{\frac{d}{dx} \frac{1}{t}}\]

OpenStudy (lgbasallote):

change that x to t

OpenStudy (turingtest):

lol

OpenStudy (anonymous):

\[\lim_{\diamondsuit\to 0^+}\frac{\ln(\diamondsuit)}{\frac{1}{\diamondsuit}}\]

OpenStudy (zarkon):

nice

OpenStudy (anonymous):

your picture here \(\diamondsuit\)

OpenStudy (lgbasallote):

\[\large\lim_{\heartsuit \rightarrow 0^+} \frac{\ln (\heartsuit)}{\frac{1}{\heartsuit}}\]

OpenStudy (lgbasallote):

@davidesm you there? lol

OpenStudy (anonymous):

I apologize, but how did you get from the original question to lnt/1/t?

OpenStudy (lgbasallote):

look at the 6th post,,maybe that'll help

OpenStudy (anonymous):

multiplying by \(t\) is the same as dividing by the reciprocal of \(t\) just like multiplying by 2 is the same as dividing my \(\frac{1}{2}\) it is just an algebra trick to turn it in to \(\frac{-\infty}{\infty}\) instead of \(0\times (-\infty)\)